### v1

# Electronic reconstruction and interface engineering of emergent spin fluctuations in compressively strained La3Ni2O7 on SrLaAlO4(001)

Abstract
Motivated by the recent observation of ambient-pressure superconductivity with K in La3Ni2O7 on SrLaAlO4(001), we explore the structural and electronic properties as well as the spin-spin correlation function of this bilayer nickelate system by using density functional theory including a Coulomb repulsion term. We find that the compressive strain exerted by this substrate leads to an unconventional occupation of the antibonding Ni states, distinct from the superconducting bulk compound under pressure. While pure strain effects rather modestly enhance the dynamical spin susceptibility via Fermi surface nesting, investigation of a reconstructed interface composition as observed in transmission electron microscopy uncovers a strong amplification of the spin fluctuations. These results provide insights into the emergence of superconductivity in strained La3Ni2O7, suggest a possible key role of the interface, and highlight fundamental differences from the hydrostatic pressure scenario.
I Introduction
The discovery of superconductivity with K in pressurized La3Ni2O7 [1, 2, 3] has identified bilayer Ruddlesden-Popper nickelates as a fascinating new addition to the growing family of superconducting nickelates. This breakthrough has sparked significant theoretical and experimental interest [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Unlike infinite-layer nickelates, where superconductivity is limited so far to thin films [40, 41, 42, 43, 44, 45, 46, 47, 48], La3Ni2O7 exhibits superconductivity in bulk crystals. The pairing mechanism has been linked to a pressure-driven Fermi surface reconstruction [1, 4, 5, 6, 16] and an associated orthorhombic-to-tetragonal structural transition [24, 30], which suppresses octahedral rotations [1, 24, 16, 30]. Depending on the balance of the involved Ni orbitals, it may correspond to -wave or symmetry [7, 16, 21, 49].
Recently, epitaxial strain has been proposed by density functional theory as a promising route to induce ambient-pressure superconductivity in La3Ni2O7 [50, 51, 52]. Enhanced spin fluctuations have been predicted particularly for tensile strain in La3Ni2O7/SrTiO3(001) [50]. Intriguingly, independent experiments have since reported the observation of ambient-pressure superconductivity with a transition onset of up to K in capped bilayer nickelate thin films of 4 to 6 bilayers grown on SrLaAlO4(001) (SLAO) [53, 54, 55], after previous attempts did not detect a superconducting transition for compressive or tensile strain [56]. This necessitates the identification of potential differences to the physics of bilayer nickelates under high pressure and how they relate to the pairing mechanism.
Here we explore the structural and electronic properties of La3Ni2O7/SrLaAlO4(001) by using density functional theory including a Coulomb repulsion term, with an explicit treatment of the interface. In addition to the ideal interface stacking, we consider a reconstructed interface composition that has recently been identified in transmission electron microscopy (TEM) [53]. Structural analysis reveals close agreement with recent x-ray diffraction (XRD) and TEM data for La3Ni2O7 on different prominent substrates [57] and emphasizes the importance of the interfacial reconstruction. We observe finite octahedral tilts throughout the bilayer nickelate, which surprisingly are quenched at the reconstructed interface due to geometric constraints. We find that compressive strain in the presence of a substrate induces an electronic reconstruction in La3Ni2O7, characterized by an unconventional partial occupation of the antibonding Ni states. Interestingly, while strain alone leads to a rather modest enhancement of the dynamical spin susceptibility as a result of Fermi surface nesting, the reconstructed interface composition substantially amplifies the spin fluctuations. A detailed disentanglement of the scattering channels reveals that the amplified fluctuations are driven by the emergent interfacial Ni states.
These results suggest that the enhanced pairing in compressively strained La3Ni2O7 arises from a different set of electronic excitations than the flat-band driven superconductivity in the bulk compound under hydrostatic pressure. The pronounced sensitivity to the interface geometry highlights the potential of interface engineering in shaping the electronic properties and spin fluctuations in bilayer nickelates, opening new avenues for manipulating their superconducting phase.
II Results
II.1 Interface geometry and electronic reconstruction in La3Ni2O7/SrLaAlO4(001)
Figure 1 shows the optimized geometry and layer-resolved density of states of epitaxial La3Ni2O7 on SLAO(001). We descibe the system by using symmetric supercells that contain two equivalent interfaces (see Methods). The substrate features a lattice constant of [53, 57] and exerts compressive strain on the bilayer nickelate, which has a pseudocubic lattice parameter of . We compare two representative interface geometries: the ideal interface, which represents a natural continuation of the two constituent structures [Fig. 1(a)], and a reconstructed geometry as suggested by recent TEM imaging [53]. The latter consists in substituting Al for Ni ions in layer L1, while the adjacent sites adopt a mixed Sr0.5La0.5 configuration instead of pure La ions, which we treat explicitly here [Fig. 1(b)].
For the ideal interface geometry [Fig. 1(a)], the vertical -site distances in the nickelate are enhanced to - due to compressive strain, compared to the bulk value of , and exhibit only minor variations. In contrast, for the reconstructed interface [Fig. 1(b)], the -site distance across layer L2 is notably reduced to due to the proximity of the AlO2 layer L1. We observe an expansion of the bilayer separation to in both systems, relative to the bulk value of .
We identify finite octahedral tilts throughout the nickelate region in both optimized geometries (Fig. 1). This contrasts with their assumed suppression in superconducting bulk La3Ni2O7 under high pressure [1, 24, 16, 30]. Interestingly, a notable exception occurs in layer L2 near the reconstructed interface, where the NiO6 octahedral tilts are quenched. This phenomenon is likely driven by the direct connectivity to the rigid, untilted AlO6 octahedra in layer L1, which impose a geometric constraint on the adjacent nickelate layer.
Along the [001] direction, each bilayer of the 327 Ruddlesden-Popper nickelate consists of three (LaO)1+ and two (NiO2)1.5- layers, yielding a nominal Ni valence. The SLAO substrate, on the other hand, is a band insulator with a formal Al3+ valence state and adopts a 214 Ruddlesden-Popper structure, consisting of single (AlO2)1- layers confined in two (Sr0.5La0.5O)0.5+ layers. At the ideal interface, the nickelate-substrate separation amounts to , while it is contracted to at the reconstructed interface (Fig. 1). This reduction can be directly attributed to the lowered formal polarity of the O layers involved. For reference, La3Ni2O7 on non-polar SrTiO3(001) presents a similar interfacial -site distance of [50].
The layer-resolved densities of states in Fig. 1 show the relative band alignment of the constituent oxides. We find in both systems that the Fermi energy is clearly located in the substrate band gap. Therefore, the Fermi surfaces shown in Fig. 2(d) are exclusively formed by the Ni states. The interface reconstruction reduces the energy difference between the Fermi level and the valence band maximum of the substrate from to eV. In contrast to La3Ni2O7/LaAlO3(001), where the polar discontinuity at the interface leads to an electron doping of the nickelate superconductor [50], we do not observe interfacial charge transfer for the present systems.
Figure 2 shows the projected momentum-resolved density of states, , where denotes Ni manifolds at different sites . In general, compressive strain drives a significant charge transfer from the Ni reservoir to the Ni states, since the concomitant vertical expansion of the nickelate region tends to lower the energy of the Ni orbital relative to Ni . Intriguingly, this allows to control the Ni orbital polarization over a considerably larger interval than hydrostatic pressure [50]. Here, we find that this mechanism leads to an unconventional occupation of the antibonding Ni states around the point [Fig 2(b)], which are empty in the bulk [Fig 2(a)], even at finite pressure. Their occupation is fairly uniform among the different layers, signaling the absence of electrostatic doping. Simultaneously, the bonding Ni states are entirely filled and their flat-band-like maximum is lowered to eV in all layers.
Importantly, this picture applies to the system with reconstructed interface as well [Fig 2(c)]. Due to the Al substitution in layer L1, the adjacent layer L2 presents a single Ni band instead of a bonding-antibonding pair [Fig 2(c)]. It is more strongly filled than the other Ni bands in either the ideal or the reconstructed system and provides a substantial contribution to the total density of states at the Fermi energy [Fig. 1(b)], which plays a central role in the spin susceptibility enhancement (see below). Analysis of the layer polarities leaves the NiO2 layer L2 in a formal state, leading to an altered local Ni3+ valence similar to bulk LaNiO3. Interestingly, this interfacial NiO2 layer is embedded in a highly anisotropic chemical environment, which induces a considerable Ni orbital polarization. This effect is even more pronounced than in LaNiO3/LaAlO3(001) superlattices, which have been widely explored in the context of Fermi surface engineering and nickelate superconductivity [58, 59].
The Fermi surface of La3Ni2O7 is known to consist of two characteristic sheets: an sheet of predominantly Ni character and a sheet with some admixture of Ni [4, 5, 6, 27]. Under hydrostatic pressure [1, 4, 5, 6, 7, 21, 16] or tensile strain as imposed by SrTiO3(001) [50], the system undergoes a topological transition, leading to the emergence of a hole pocket due to a metallization of the bonding Ni states. In stark contrast, our simulations for compressively strained La3Ni2O7/SLAO(001) reveal a distinct Fermi surface composition, reflecting the differences in electronic reconstruction [Fig. 2(d)]. The sheet exhibits an enhanced Ni contribution compared to the bulk, while the sheet appears more rectangular and features a significantly promoted Ni character. Additional electron pockets emerge near the point, originating from the partial occupation of the antibonding Ni states (). For the reconstructed interface composition, we find a notable splitting of the sheet. Moreover, a larger electron pocket around the point can be observed, predominantly associated with Ni states localized in layer L2, and a smaller pocket that stems from layer L3 [Fig. 2(c,d)]. These results highlight the effect of interface reconstructions on the electronic structure and suggest their critical role in the modulation of the superconducting properties. We will corroborate this aspect by calculating the dynamical spin susceptibility below.
II.2 Structural properties of bilayer nickelates on different substrates
The availability of high-quality x-ray diffraction and TEM data on strained epitaxial thin films [57], specifically for the evolution of the vertical lattice parameter of the bilayer nickelate, offers a unique opportunity to systematically compare experimental observations with the optimized structures reported here and in previous work [50]. Extracting an effective nickelate lattice parameter from the optimized structures can be performed in different ways, which yield slightly different values. In Fig. 3, we present obtained by using two distinct methods: (i) The outer distance (filled diamonds) is defined as the distance between the two interfaces. This corresponds to of an orthorhombic La3Ni2O7 reference cell. (ii) The inner distance (empty diamonds) corresponds to and captures more specifically the internal structural response of the nickelate layer.
Interestingly, we find that the outer distance follows approximately a function (dashed black line), a behavior observed also in LaNiO3/LaAlO3(001) superlattices [61]. For La3Ni2O7/SLAO(001), it yields (ideal interface) and (reconstructed interface). Both values clearly exceed the bulk reference of [24], indicating a significant strain-induced vertical expansion of the nickelate. The inner distance, which generally renders slightly larger values, yields (ideal interface) and (reconstructed interface). We observe that the reconstructed interface geometry reduces considerably, bringing the theoretical predictions for La3Ni2O7/SLAO(001) into close agreement with the experimental TEM and XRD results (Fig. 3). For La3Ni2O7/SrTiO3(001) (LNO/STO), we obtain very good agreement between theoretical and experimental values as well. For La3Ni2O7/LaAlO3(001) (LNO/LAO), the inner distance aligns closely with the experimental data, whereas the outer distance is notably higher, reflecting the complex interplay of interfacial charge transfer and structural distortions in this system.
Overall, our DFT predictions agree within a margin with the experimental lattice constants over a wide range of strain values from to , despite systematic uncertainties in the theoretical and experimental data. This comparison underscores the impact of interface effects on the structural parameters. The values compiled in Fig. 3 will be a helpful reference for future experimental studies.
II.3 Spin susceptibility in La3Ni2O7/SrLaAlO4(001)
The superconducting pairing is determined by the dynamical spin susceptibility , which we compare in Fig. 4 considering a Kanamori-type interaction vertex with and eV in the random phase approximation (RPA). For completeness, we also show the bare Lindhard susceptibility .
The spin susceptibility of bulk La3Ni2O7 at ambient pressure is dominated by intraorbital scattering [50]. Direct comparison reveals that La3Ni2O7/SLAO(001) exhibits a moderately enhanced magnitude and sharpness of the characteristic susceptibility peaks along the direction. The difference can be traced back to the intraorbital and interorbital - channels (not shown).
Intriguingly, we find a very strong enhancement of the susceptibility for the La3Ni2O7/SLAO(001) system if we consider the reconstructed interface geometry observed in TEM. These changes are not only quantitative, but also qualitative, as they are driven by considerably increased intraorbital scattering and further promoted by contributions from the interorbital - channel (Fig. 4, bottom row). Both clearly surpass the channel.
We identify that this phenomenon originates predominantly from the emergent Ni states localized near the interface, which contribute strongly to the density of states at the Fermi level and exhibit favorable Fermi surface nesting properties (Figs. 1 and 2). Together, these effects result in the strong amplification of the spin fluctuations. Considering that recent experiments reporting ambient-pressure superconductivity in strained systems are performed on capped bilayer nickelate films of 4 to 6 bilayers thickness [53, 54, 55], these findings suggest that interfacial effects may play a key role in the emergence of superconductivity in La3Ni2O7 on SLAO(001).
III Discussion
We investigated the effect of compressive strain and interface reconstructions in La3Ni2O7 on SrLaAlO4(001) by performing first-principles simulations including a Coulomb repulsion term, treating the interface explicitly.
Comparison of the structural optimization results, specifically the evolution of the vertical lattice parameter of the bilayer nickelate on different technologically relevant substrates, with recent x-ray diffraction and transmission electron microscopy measurements on strained epitaxial thin films revealed close agreement and highlighted the key role of the reconstructed interface composition. We observed finite octahedral tilts throughout the bilayer nickelate. Intriguingly, these tilts are suppressed at the reconstructed interface, accompanied by significantly enhanced spin fluctuations. This finding provides an important contribution to the ongoing debate on whether superconductivity in bilayer nickelates requires a structural transition to the high-symmetry phase or can arise from purely electronic effects [24, 50].
We found that compressive strain drives an electronic reconstruction in the bilayer nickelate, characterized by an unconventional occupation of the antibonding Ni states with simultaneous absence of interfacial charge transfer. This scenario contrasts sharply with the notion of metallized bonding Ni states, considered to be critical for the superconducting state in pressurized bulk La3Ni2O7. The concomitant substantial modifications of the Fermi surface and amplified spin fluctuations, especially for the reconstructed interface due to the emergent Ni states, point to an electron pairing that deviates fundamentally from the flat-band physics dominant under hydrostatic pressure.
The pronounced influence of epitaxial strain and interface effects underscores the potential of strain and interface engineering to tailor the electronic structure and optimize the superconducting properties in bilayer nickelates at ambient pressure, circumventing the challenges associated with high-pressure synthesis and characterization and offering a pathway towards realizing their full technological potential.
Acknowledgements.
This work was supported by the National Science Foundation, Grant No. NSF-DMR-2118718.IV Methods
IV.1 Density functional theory calculations
We performed first-principles calculations in the framework of density functional theory [62] (DFT) as implemented in the Quantum Espresso code [63]. The generalized gradient approximation was used for the exchange and correlation functional as parametrized by Perdew, Burke, and Ernzerhof [64] in conjunction with ultrasoft pseudopotentials [65], as successfully employed in previous work [66, 67, 61]. Static correlation effects were considered within the DFT formalism [68], employing eV on the Ni and Ti states [69, 70, 71, 1, 16, 50]. Octahedral rotations are fully accounted for by using supercells with two transition metal sites per layer, setting to the substrate lattice parameter of [53, 57] and accurately optimizing as well as the ionic positions. The symmetric supercells contain two equivalent interfaces, 3 bilayers of La3Ni2O7, and 5 layers of substrate, corresponding to 142 atoms (Fig. 1). Wave functions and density were expanded into plane waves up to cutoff energies of and Ry, respectively. The Brillouin zone was sampled by using a Monkhorst-Pack -point grid [72] and mRy Methfessel-Paxton smearing [73]. The ionic positions were accurately optimized, reducing ionic forces below mRya.u. Subsequently, Fermi surfaces and densities of states were obtained by using -point grids. Reference calculations for ambient-pressure bulk La3Ni2O7 () were performed by using 48-atom supercells [24, 34] with and -point grids.
IV.2 Tight-binding Hamiltonians in a basis of maximally localized Wannier functions
We constructed tight-binding Hamiltonians describing the physics of the Ni states, which are the active states near the Fermi level, in a basis of maximally localized Wannier functions for all bilayer nickelate systems considered in this work. The tight-binding models describe 12 (8) distinct Ni sites and thus 24 (16) Ni orbitals for the system with ideal (reconstructed) interface, which can be further reduced by exploiting symmetries. The output files of wannier90 [74], which can be directly imported into different post-processing programs, are conveniently provided at https://github.com/henniggroup.
IV.3 Spin susceptibility calculations
Using these tight-binding Hamiltonians, we computed the bare generalized susceptibility at K () leveraging the Toolbox for Research on Interacting Quantum Systems-Two-Particle Response Function (TRIQS-TPRF) framework [75] by Fourier transform of
from non-interacting single-particle Green’s functions defined on a mesh of Matsubara frequencies
where is the matrix-valued dispersion relation. The indices label the different orbitals ( and ) at the different Ni sites. Note that we employ the full tight-binding Hamiltonian based on the supercell here instead of a simplified model, describing explicitly all Ni sites to accurately treat effects of octahedral rotations and electrostatic doping. Subsequently, we consider a Kanamori-type interaction vertex during the calculation of in the random phase approximation (RPA),
with
where and . Notably, characterizes electronic correlations involving orbitals at the same Ni site, vanishing if any indices correspond to distinct sites. Finally, we implemented a strategy to unfold the result from the to the full Brillouin zone by including structural phase factors and computed the ’physical’ susceptibility by evaluating the trace:
Leveraging this methodology, we are able to fully consider the effects of octahedral rotations and electrostatic doping that necessitate the use of large supercells, while displaying without backfolding for direct comparison with future experiments.